The larger of two numbers is 1 more than twice the smaller. The sum of the numbers is 20 less than three times the larger. Find the number
I figured out the two equations which are
x = smaller
y = larger

y = 2x+1 ; x+y = 3y-20
I found the equations but I don't know how to solve it, to get the numbers of x and y.

Answer by ewatrrr(24785)


Hi
y = 2x+1
x+y = 3y-20 |good work!
x + (2x+1) = 3(2x+1) - 20 |substituting for y
solving for x
3x + 1 = 6x + 3 - 20
18 = 3x
x = 6 and y = 13 (2x+1)
CHECKING our Answer***
y = 2x+1
13 = 2*6 + 1